### v1

# Complexity and curvature of pairs of Burch modules and ideals

Abstract.
The complexity and curvature of a module were first introduced by Avramov to distinguish modules of infinite homological dimension. Later, Avramov–Buchweitz extended the notion of complexity from a single module to that of pairs of modules, which measures the polynomial growth rate of the minimal number of generators of their Ext modules. Dao studied a similar notion of Tor-complexity. Recently, Dey–Ghosh–Saha initiated the study of Ext and Tor curvature of a pair of modules, which measure the exponential growth rates of the corresponding Ext and Tor, respectively. On the other hand, the concept of Burch ideals was introduced by Dao–Kobayashi–Takahashi, motivated by classical work of Burch. This class includes several large and well-studied families of ideals in a Noetherian local ring $(R,\mathfrak{m},k)$. For example, every nonzero ideal of the form $\mathfrak{a}\mathfrak{m}$ (e.g., $\mathfrak{m}^{n}$ for $n\geqslant 1$), and under mild conditions every integrally closed ideal $I$ with $\operatorname{depth}(R/I)=0$, is a Burch ideal. Suppose $I$ and $J$ are Burch ideals such that $I$ is $\mathfrak{m}$-primary. One of the main results of this article is $\operatorname{cx}_{R}(I,J)=\operatorname{tcx}_{R}(I,J)=\operatorname{cx}_{R}(k)$. Moreover, we show that $R$ is complete intersection if and only if $\operatorname{cx}_{R}(I,J)$ or $\operatorname{tcx}_{R}(I,J)$ is finite if and only if $\operatorname{curv}_{R}(I,J)$ or $\operatorname{tcurv}_{R}(I,J)$ is at most $1$. We deduce these results from the corresponding more general results on Burch modules (in the sense of Dey-Kobayashi).
Key words and phrases:
Complete intersection rings; Burch ideals and modules; Ext and Tor; Complexity and curvature
2020 Mathematics Subject Classification:
13D07, 13D05, 13C13, 13H10
1. Introduction
Throughout the article, unless specified otherwise, we assume $(R,\mathfrak{m},k)$ to be a commutative Noetherian local ring. All modules over $R$ are assumed to be finitely generated.
The notions of complexity and curvature of a module were originally introduced by Avramov in [Avr89a], [Avr89b] and [Avr96], see Definitions 2.1 and 2.3. These invariants distinguish modules of infinite projective and injective dimensions. Later, in [AB00], Avramov and Buchweitz extended the notion of complexity to pairs of modules $(M,N)$, which measures the polynomial growth rate of minimal number of generators $\mu(\operatorname{Ext}^{n}_{R}(M,N))$ of the Ext modules as $n$ increases. Dao introduced the analogous concept of Tor-complexity in [Dao07]. This measures the polynomial growth rate of $\mu(\operatorname{Tor}_{n}^{R}(M,N))$. Influenced by these, in [DGS24], Dey, Ghosh and Saha developed the concept of Ext and Tor curvatures for a pair of modules; these measure the exponential growth rate of Ext and Tor modules, respectively, see Definition 2.4. Additionally, in [DGS24, Thms. A and B, Cors. 5.5 and 6.6], Dey, Ghosh and Saha established various inequalities between (Ext and Tor) complexity and curvature of pairs of certain CM (Cohen-Macaulay) modules and that of a single module. Along with other applications, they also provided a number of criteria for a local ring to be complete intersection (including hypersurface and regular local ring) via the complexity and curvature of pairs of CM modules satisfying certain inequalities involving multiplicity, see [DGS24, Sec. 7]. In the present article, we study the (Ext and Tor) complexity and curvature of pairs of Burch modules, and establish various (in)equalities. Along the way, we obtain a number of new characterizations of complete intersection local rings (including hypersurface rings). Recall that $R$ is said to be a complete intersection of codimension $c$ if its completion $\widehat{R}=Q/({\bf f})$, where $(Q,\mathfrak{n})$ is a regular local ring, and ${\bf f}\in\mathfrak{n}^{2}$ is a $Q$-regular sequence of length $c$. If $R$ is complete intersection of codimension at most $1$, then $R$ is called a hypersurface.
The class of Burch ideals and modules is an important class of ideals and modules studied in commutative algebra. In [DKT20], Dao, Kobayashi and Takahashi introduced the notion of Burch ideals motivated by some striking results of Burch [burch, p. 946, Thm. 5.(ii), and p. 947, Cor. 1.(ii)]. Later, Dey and Kobayashi extended this concept, and introduced the notion of Burch submodules in [DK23]. An $R$-module $M$ is said to be a Burch submodule of an $R$-module $L$ if $\mathfrak{m}(M:_{L}\mathfrak{m})\neq\mathfrak{m}M$. A Burch submodule of $R$ is called a Burch ideal. An $R$-module $M$ is called Burch if it is isomorphic to a Burch submodule of some $R$-module. There are several well-studied classes of examples of Burch (sub)modules and ideals. Here we list some interesting classes.
Example 1.1.
  1. (1)
    For a submodule $M$ of an $R$-module $L$, if $\mathfrak{m}M\neq 0$, then $\mathfrak{m}M$ is a Burch submodule of $L$, see [DK23, Ex. 3.4]. Moreover, a non-zero ideal $I$ of the form $\mathfrak{a}\mathfrak{m}$ (for some ideal $\mathfrak{a}$) is a Burch ideal. In particular, if $\mathfrak{m}^{n}\neq 0$ for some $n\geqslant 1$, then $\mathfrak{m}^{n}$ is an $\mathfrak{m}$-primary Burch ideal.
  2. (2)
    It follows from (1) and [DK23, Lem. 5.2] that Ulrich modules of dimension 1 are Burch.
  3. (3)
    If $M$ is Burch, then so is $M\oplus X$ for any $R$-module $X$.
  4. (4)
    Under some mild conditions on $R$, it is shown in [GS24, Prop. 4.6] that every integrally closed ideal $I$ such that $\operatorname{depth}(R/I)=0$ is Burch; in particular, $\mathfrak{m}$-primary integrally closed ideals are Burch ideals.
  5. (5)
    If $R$ is of prime characteristic $p$, $F$-finite, and $\operatorname{depth}(R)=1$, then the $e$-th Frobenius pushforward ${{}^{e}\!}{R}$ is Burch for all $e\gg 0$, see [CDKMS, 2.4.(vi)]. Recall that ${{}^{e}\!}{R}$ denotes the $R$-module whose underlying abelian group is $R$, and the $R$-action on ${{}^{e}\!}{R}$ is given by $r\cdot x=r^{p^{e}}x$ for $r\in R$ and $x\in{{}^{e}\!}{R}$.
Burch ideals and modules possess several nice homological properties. For instance, they are strong test modules that detect finite homological dimensions via finite vanishing of Ext or Tor ([burch, p. 946, Thm. 5.(ii)]; [DK23, Prop. 3.16]). They characterize regular, complete intersection, Gorenstein, and CM local rings in terms of the corresponding homological dimensions ([burch, p. 947, Cor. 1.(ii)]; [DK23, Cor. 3.19]; [GS24, Thms. 1.1 and 5.1.(6)]). They also satisfy the Huneke–Wiegand conjecture over Gorenstein rings [CDKMS, Thm. 1.2].
Towards understanding growth in minimal free and injective resolutions of Burch modules, it follows from a result of Avramov [Avr96, Thm. 4] that a Burch module (and hence a Burch ideal) has maximal (injective) complexity and curvature; see Proposition 2.20. As a consequence of this and [Avr96, Thm. 3], it follows that $R$ is a complete intersection if (injective) complexity of $M$ is finite, or (injective) curvature of $M$ is at most $1$, where $M$ is some Burch $R$-module. Since $\mathfrak{m}$ is always Burch (unless $R$ is a field), and $\operatorname{cx}_{R}(M)=\operatorname{tcx}_{R}(\mathfrak{m},M)$, $\operatorname{curv}_{R}(M)=\operatorname{tcurv}_{R}(\mathfrak{m},M)$, $\operatorname{inj\,cx}_{R}(M)=\operatorname{cx}_{R}(\mathfrak{m},M)$, $\operatorname{inj\,curv}_{R}(M)=\operatorname{curv}_{R}(\mathfrak{m},M)$, it is natural to ask the following questions.
Question 1.2.
Let $M$ and $N$ be Burch modules.
  1. (1)
    If $\operatorname{tcx}_{R}(M,N)$ or $\operatorname{cx}_{R}(M,N)$ is finite, then is $R$ a complete intersection?
  2. (2)
    If $\operatorname{tcurv}_{R}(M,N)\leqslant 1$ or $\operatorname{curv}_{R}(M,N)\leqslant 1$, then is $R$ a complete intersection?
Under certain annihilation conditions on $\operatorname{Ext}$ and $\operatorname{Tor}$, we establish various (in)equalities between the complexity (resp., curvature) of a pair of (Burch) modules and that of a single module; see Propositions 3.1, 3.3 and 3.5, and Theorems 3.2, 3.4, 3.6, 3.10 and 3.15. Combining Remark 3.16 with some of these results, and a key insight coming from Lemma 2.19, we obtain the following main results of this paper which provide partial affirmative answers to 1.2.
Theorem 1.3 (See Theorems 3.10 and 3.13 for stronger results).
Assume that $R$ is CM. Let $M$ and $N$ be $R$-modules such that $N$ is Burch. Suppose that either
  1. (1)
    $M$ locally has finite projective dimension on the punctured spectrum, or
  2. (2)
    $\widehat{R}$ is locally Gorenstein on the punctured spectrum ((e.g., see Remark 3.14)), and $N$ locally has finite projective dimension on the punctured spectrum.
Then $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,N)$ and $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$.
Theorem 1.4 (See Theorem 3.15 for stronger results).
Let $M$ and $N$ be $R$-modules such that $M$ is a Burch submodule of some $R$-module $X$. Assume that $\operatorname{G-dim}_{R}(X)<\infty$, or $X$ is torsionless. Then, the following hold.
  1. (1)
    If $M$ or $N$ locally has finite projective dimension on the punctured spectrum, then
    $\operatorname{cx}_{R}(N)\leqslant\operatorname{tcx}_{R}(M,N)\;\mbox{ and }\;\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$
  2. (2)
    If $M$ locally has finite projective dimension on the punctured spectrum, then
    $\operatorname{inj\,cx}_{R}(N)\leqslant\operatorname{cx}_{R}(M,N)\;\mbox{ and }\;\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$
Thus, we also characterize complete intersection local rings (including hypersurface rings) in terms of (Ext and Tor) complexity and curvature of pairs of Burch modules and ideals. Certain special cases of these results yield the following corollaries.
Corollary 1.5.
Let $I$ and $J$ be two Burch ideals of $R$ ((e.g., II and $J$ can be any of the ideals as described in Example 1.1)). Suppose that $I$ is $\mathfrak{m}$-primary. Then $\operatorname{tcx}_{R}(I,J)=\operatorname{cx}_{R}(I,J)=\operatorname{cx}_{R}(k)$. Moreover, the following conditions are equivalent:
  1. (1)
    $R$ is complete intersection ((resp., of codimension $c$)).
  2. (2)
    $\operatorname{tcx}_{R}(I,J)<\infty$ ((resp., $\operatorname{tcx}_{R}(I,J)=c<\infty$)).
  3. (3)
    $\operatorname{cx}_{R}(I,J)<\infty$ ((resp., $\operatorname{cx}_{R}(I,J)=c<\infty$)).
  4. (4)
    $\operatorname{tcurv}_{R}(I,J)\leqslant 1$.
  5. (5)
    $\operatorname{curv}_{R}(I,J)\leqslant 1$.
Consequently, $R$ is a hypersurface $\Longleftrightarrow$ $\operatorname{tcx}_{R}(I,J)\leqslant 1$ $\Longleftrightarrow$ $\operatorname{cx}_{R}(I,J)\leqslant 1$.
For $R$-modules $M$ and $N$, note that $\operatorname{tcx}_{R}(M,N)=\operatorname{tcx}_{R}(N,M)$. However, $\operatorname{cx}_{R}(M,N)$ might not be the same as $\operatorname{cx}_{R}(N,M)$. So, a natural question arises whether the conclusions of Corollary 1.5 hold if $J$ is $\mathfrak{m}$-primary instead of that $I$ is $\mathfrak{m}$-primary? The next result particularly ensures that the answer to this question is affirmative when $R$ is Gorenstein (more generally, when $R$ is CM such that $\widehat{R}$ is locally Gorenstein on the punctured spectrum, cf. Remark 3.14).
Corollary 1.6.
Let $R$ be CM such that $\widehat{R}$ is locally Gorenstein on the punctured spectrum ((e.g., the rings in Remark 3.14)). Let $I$ and $J$ be two Burch ideals of $R$ ((e.g., II and $J$ can be any of the ideals as described in Example 1.1)). Suppose that $J$ is $\mathfrak{m}$-primary. Then $\operatorname{cx}_{R}(I,J)=\operatorname{cx}_{R}(k)$. Moreover, the following conditions are equivalent:
  1. (1)
    $R$ is complete intersection ((resp., of codimension $c$)).
  2. (2)
    $\operatorname{cx}_{R}(I,J)<\infty$ ((resp., $\operatorname{cx}_{R}(I,J)=c<\infty$)).
  3. (3)
    $\operatorname{curv}_{R}(I,J)\leqslant 1$.
Consequently, $R$ is a hypersurface $\Longleftrightarrow$ $\operatorname{cx}_{R}(I,J)\leqslant 1$.
The organization of this article is as follows. In Section 2, we set up our main notions and notations, and recall the primary definitions, lemmas, and propositions for use throughout the rest of the paper. Section 3 contains the proofs of the main results of this article, including the proofs of Theorems 1.3 and 1.4, as well as Corollaries 1.5 and 1.6. Finally, in Section 4, we compute some examples that support our main results.
Acknowledgments
Souvik Dey was partly supported by the Charles University Research Center program No.UNCE/24/SCI/022 and a grant GAČR 23-05148S from the Czech Science Foundation. This work is part of the third named author’s Ph.D. thesis. She would like to thank the Government of India for financial support through the Prime Minister Research Fellowship for her Ph.D.
2. Preliminaries
In this section, we recall essential notation, definitions, and fundamental properties that are used in the rest of the article. For an $R$-module $M$, the $n$th Betti number of $M$ is defined to be $\beta_{n}^{R}(M):=\operatorname{rank}_{k}(\operatorname{Ext}^{n}_{R}(M,k))$, while the $n$th Bass number of $M$ is defined by $\mu_{R}^{n}(M):=\operatorname{rank}_{k}(\operatorname{Ext}^{n}_{R}(k,M))$. For $n\geqslant 0$, the $n$th syzygy of $M$ in a minimal free resolution is denoted by $\operatorname{\Omega}^{n}_{R}(M)$. We denote by $\mu_{R}(M)$ and $\lambda_{R}(M)$ the minimal number of generators and the length of $M$, respectively. Set $\operatorname{ann}_{R}(M):=\{r\in R:rx=0\mbox{ for all }x\in M\}$. Let $\widehat{(-)}$ denote the $\mathfrak{m}$-adic completion functor. For two $R$-modules $M$ and $N$, the notation $\operatorname{Ext}^{\gg 0}_{R}(M,N)$ (resp., $\operatorname{Tor}_{\gg 0}^{R}(M,N)$) is used to denote all $R$-modules $\operatorname{Ext}^{n}_{R}(M,N)$ (resp., $\operatorname{Tor}_{n}^{R}(M,N)$) for sufficiently large $n$.
Definition 2.1.
Let $\{a_{n}\}$ be a sequence of non-negative real numbers.
  1. (1)
    The complexity of $\{a_{n}\}$, denoted by $\operatorname{cx}\{a_{n}\}$, is defined to be the smallest integer $b\geqslant 0$ such that $a_{n}\leqslant\alpha n^{b-1}$ for all $n\gg 0$ and for some constant $\alpha>0$. If no such $b$ exists, then $\operatorname{cx}\{a_{n}\}:=\infty$.
  2. (2)
    The curvature of $\{a_{n}\}$ is defined by $\operatorname{curv}\{a_{n}\}:=\displaystyle\limsup_{n\to\infty}\sqrt[n]{a_{n}}$.
Lemma 2.2.
[DGS24, Lem. 2.7 and Rmk. 2.2.(1)] For sequences $\{x_{n}\}$ and $\{y_{n}\}$ of non-negative integers,
  1. (1)
    $\operatorname{cx}(\{x_{n}+y_{n}\})=\sup\{\operatorname{cx}(\{x_{n}\}),\operatorname{cx}(\{y_{n}\})\}$.
  2. (2)
    $\operatorname{curv}(\{x_{n}+y_{n}\})=\sup\{\operatorname{curv}(\{x_{n}\}),\operatorname{curv}(\{y_{n}\})\}$.
The complexity of a module was introduced by Avramov in [Avr89a, Def. 3.1] and [Avr89b, Def. 1.1], while the injective complexity, referred to as plexity, was defined in [Avr89b, Def. 5.1]. The study of (injective) curvature was initiated in [Avr96, Sec. 1].
Definition 2.3 (Avramov).
Let $M$ be an $R$-module.
  1. (1)
    The complexity and injective complexity of $M$ are defined to be cx_R(M)=cx{β^R_n(M)} and inj cx_R(M)=cx{μ_R^n(M)}.
  2. (2)
    The curvature and injective curvature of $M$ are defined by curv_R(M)= curv{β^R_n(M)}  and  inj curv_R(M)= curv{μ_R^n(M)}.
The concepts of complexity and curvature were extended from a single module to a pair of modules by Avramov-Buchweitz, Dao and Dey-Ghosh-Saha.
Definition 2.4.
Let $M$ and $N$ be two $R$-modules.
  1. (1)
    [AB00, p. 286] The complexity (or Ext-complexity) of $M$ and $N$ is defined to be cx_R(M,N):=cx{μ_R(Ext_R^n(M,N))}_n⩾0.
  2. (2)
    [DGS24, Def. 3.2] The curvature (or Ext-curvature) of $M$ and $N$ is defined by curv_R(M,N):=curv{μ_R(Ext_R^n(M,N))}_n ⩾0.
  3. (3)
    [Dao07, p. 4] The Tor-complexity of $M$ and $N$ is defined to be tcx_R(M,N):=cx{μ_R(Tor^R_n(M,N))}_n⩾0.
  4. (4)
    [DGS24, Def. 3.3] The Tor-curvature of $M$ and $N$ is defined by tcurv_R(M,N):=curv{μ_R(Tor^R_n(M,N))}_n ⩾0.
Remark 2.5.
The Tor-curvature is also defined in [Sega, 3.8] in terms of length of Tor modules. However, the definitions in terms of $\mu$ and in terms of $\lambda$ are equivalent when $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$; see [DGS24, Lem. 3.4].
Remark 2.6.
For $R$-modules $M$ and $N$, the following hold:
  1. (1)
    $\operatorname{cx}_{R}(M)=\operatorname{tcx}_{R}(M,k)=\operatorname{cx}_{R}(M,k)$ and $\operatorname{inj\,cx}_{R}(M)=\operatorname{cx}_{R}(k,M)$.
  2. (2)
    $\operatorname{curv}_{R}(M)=\operatorname{tcurv}_{R}(M,k)=\operatorname{curv}_{R}(M,k)$ and $\operatorname{inj\,curv}_{R}(M)=\operatorname{curv}_{R}(k,M)$.
  3. (3)
    [DGS24, Rmk. 2.3 and 3.8.(3)]
    1. (a)
      $\operatorname{cx}_{R}(M,N)<\infty$ $\Longrightarrow$ $\operatorname{curv}_{R}(M,N)\leqslant 1$ $\Longleftrightarrow$ $\operatorname{curv}_{R}(M,N)=0$ or $1$.
    2. (b)
      $\operatorname{tcx}_{R}(M,N)<\infty$ $\Longrightarrow$ $\operatorname{tcurv}_{R}(M,N)\leqslant 1$ $\Longleftrightarrow$ $\operatorname{tcurv}_{R}(M,N)=0$ or $1$.
    3. (c)
      $\operatorname{cx}_{R}(M,N)=0$ $\Longleftrightarrow$ $\operatorname{curv}_{R}(M,N)=0$ $\Longleftrightarrow$ $\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$.
    4. (d)
      $\operatorname{tcx}_{R}(M,N)=0$ $\Longleftrightarrow$ $\operatorname{tcurv}_{R}(M,N)=0$ $\Longleftrightarrow$ $\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$.
Among all the $R$-modules, the residue field of $R$ has maximal (injective) complexity and curvature. These also characterize complete intersection local rings.
Proposition 2.7.
  1. (1)
    [Avr96, Prop. 2] For any $R$-module $M$, the following hold.
    1. (i)
      $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(k)$ and $\operatorname{inj\,cx}_{R}(M)\leqslant\operatorname{inj\,cx}_{R}(k)=\operatorname{cx}_{R}(k)$.
    2. (ii)
      $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(k)$ and $\operatorname{inj\,curv}_{R}(M)\leqslant\operatorname{inj\,curv}_{R}(k)=\operatorname{curv}_{R}(k)$.
  2. (2)
    [Avr96, Thm. 3] The following conditions are equivalent:
    1. (i)
      $R$ is complete intersection ((resp., of codimension $c$)).
    2. (ii)
      $\operatorname{cx}_{R}(k)<\infty$ ((resp., $\operatorname{cx}_{R}(k)=c<\infty$)).
    3. (iii)
      $\operatorname{curv}_{R}(k)\leqslant 1$.
The (injective) complexity and curvature are preserved under taking quotient by a regular element of both the ring and the module.
Lemma 2.8.
Let $M$ be an $R$-module. Let $x$ be an element of $R$ which is regular on both $R$ and $M$. Then
  1. (1)
    $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R/xR}(M/xM)$ and $\operatorname{curv}_{R}(M)=\operatorname{curv}_{R/xR}(M/xM)$.
  2. (2)
    $\operatorname{inj\,cx}_{R}(M)=\operatorname{inj\,cx}_{R/xR}(M/xM)$ and $\operatorname{inj\,curv}_{R}(M)=\operatorname{inj\,curv}_{R/xR}(M/xM)$.
Proof.
In view of [Mat89, p. 140, Lem. 2], $\operatorname{Ext}_{R}^{n}(M,k)\cong\operatorname{Ext}_{R/xR}^{n}(M/xM,k)$ and $\operatorname{Ext}_{R}^{n+1}(k,M)\cong\operatorname{Ext}_{R/xR}^{n}(k,M/xM)$ for all $n\geqslant 0$. So $\beta_{n}^{R}(M)=\beta_{n}^{R/xR}(M/xM)$ and $\mu^{n+1}_{R}(M)=\mu^{n}_{R/xR}(M/xM)$ for all $n\geqslant 0$. Hence the desired equalities follow. ∎
Under certain conditions, (Ext and Tor) complexity of a pair of modules can be bounded above by complexity of each of the modules.
Proposition 2.9.
Let $R$ be complete intersection, and $M$, NN be two $R$-modules. Then the following hold.
  1. (1)
    $\operatorname{cx}_{R}(M)=\operatorname{inj\,cx}_{R}(M)$, and $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}<\infty$ and $\operatorname{curv}_{R}(M,N)\leqslant 1$.
  2. (2)
    Suppose that $\mathfrak{m}^{h}\operatorname{Tor}_{\gg 0}^{R}(M,N)=0$ for some $h\geqslant 0$. Then $\operatorname{tcx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}<\infty$ and $\operatorname{tcurv}_{R}(M,N)\leqslant 1$.
Proof.
(1) It is shown in [AB00, Thm. II] that $\operatorname{cx}_{R}(M)=\operatorname{inj\,cx}_{R}(M)$ and $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}$. By Proposition 2.7.(1), $\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(k)<\infty$. Thus $\operatorname{cx}_{R}(M,N)<\infty$. Hence, in view of Remark 2.6.(3), $\operatorname{curv}_{R}(M,N)\leqslant 1$.
(2) The inequalities for (Tor) complexities are shown in [Dao07, Cor. 5.7] and [DGS24, Prop. 7.9.(1)]. Hence, by Remark 2.6.(3), $\operatorname{tcurv}_{R}(M,N)\leqslant 1$. ∎
Both complexity and curvature behave well with finite direct sum of modules.
Lemma 2.10.
Let $M$, NN, $M_{i}$ and $N_{i}$ be $R$-modules for $i=1,2$. Then
  1. (1)
    $\operatorname{tcx}_{R}(M_{1}\oplus M_{2},N)=\sup\{\operatorname{tcx}_{R}(M_{1},N),\operatorname{tcx}_{R}(M_{2},N)\}$.
  2. (2)
    $\operatorname{tcurv}_{R}(M_{1}\oplus M_{2},N)=\sup\{\operatorname{tcurv}_{R}(M_{1},N),\operatorname{tcurv}_{R}(M_{2},N)\}$.
  3. (3)
    $\operatorname{cx}_{R}(M_{1}\oplus M_{2},N)=\sup\{\operatorname{cx}_{R}(M_{1},N),\operatorname{cx}_{R}(M_{2},N)\}$.
  4. (4)
    $\operatorname{curv}_{R}(M_{1}\oplus M_{2},N)=\sup\{\operatorname{curv}_{R}(M_{1},N),\operatorname{curv}_{R}(M_{2},N)\}$.
  5. (5)
    $\operatorname{cx}_{R}(M,N_{1}\oplus N_{2})=\sup\{\operatorname{cx}_{R}(M,N_{1}),\operatorname{cx}_{R}(M,N_{2})\}$.
  6. (6)
    $\operatorname{curv}_{R}(M,N_{1}\oplus N_{2})=\sup\{\operatorname{curv}_{R}(M,N_{1}),\operatorname{curv}_{R}(M,N_{2})\}$.
Proof.
Note that for any two $R$-modules $L_{1}$ and $L_{2}$, one has that $\mu_{R}(L_{1}\oplus L_{2})=\mu_{R}(L_{1})+\mu_{R}(L_{2})$. Hence, since $\operatorname{Tor}^{R}_{n}(M_{1}\oplus M_{2},N)\cong\operatorname{Tor}^{R}_{n}(M_{1},N)\oplus\operatorname{Tor}^{R}_{n}(M_{2},N)$ for all $n$, the proofs of (1) and (2) are direct consequences of Lemma 2.2. Similarly, considering Ext modules, one obtains the other equalities. ∎
2.11.
Let $X\to Y\to Z$ be an exact sequence of $R$-modules. Then, $(\operatorname{ann}_{R}X)(\operatorname{ann}_{R}Z)\subseteq\operatorname{ann}_{R}Y$. Moreover, if all $X$, $Y$ and $Z$ have finite lengths, then $\lambda_{R}(Y)\leqslant\lambda_{R}(X)+\lambda_{R}(Z)$.
Lemma 2.12.
Let $L,M_{1},M_{2},M_{3}$ and $N$ be $R$-modules such that there is a short exact sequence $0\to M_{1}\to M_{2}\to M_{3}\to 0$. Let $s,t,u\in\{1,2,3\}$ be distinct ((e.g., $s=2$, $t=1$ and $u=3$)).
  1. (1)
    Suppose that $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M_{s},N)=0$ and $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M_{t},N)=0$ for some $h>0$. Then, $\mathfrak{m}^{2h}\operatorname{Tor}^{R}_{\gg 0}(M_{u},N)=0$. Moreover, $\operatorname{tcx}_{R}(M_{u},N)\leqslant\sup\{\operatorname{tcx}_{R}(M_{s},N),\operatorname{tcx}_{R}(M_{t},N)\}$ and $\operatorname{tcurv}_{R}(M_{u},N)\leqslant\sup\{\operatorname{tcurv}_{R}(M_{s},N),\operatorname{tcurv}_{R}(M_{t},N)\}$.
  2. (2)
    Suppose that $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(M_{s},N)=0$ and $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(M_{t},N)=0$ for some $h>0$. Then, $\mathfrak{m}^{2h}\operatorname{Ext}^{\gg 0}_{R}(M_{u},N)=0$. Moreover, $\operatorname{cx}_{R}(M_{u},N)\leqslant\sup\{\operatorname{cx}_{R}(M_{s},N),\operatorname{cx}_{R}(M_{t},N)\}$ and $\operatorname{curv}_{R}(M_{u},N)\leqslant\sup\{\operatorname{curv}_{R}(M_{s},N),\operatorname{curv}_{R}(M_{t},N)\}$.
  3. (3)
    Suppose that $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(L,M_{s})=0$ and $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(L,M_{t})=0$ for some $h>0$. Then, $\mathfrak{m}^{2h}\operatorname{Ext}^{\gg 0}_{R}(L,M_{u})=0$. Moreover, $\operatorname{cx}_{R}(L,M_{u})\leqslant\sup\{\operatorname{cx}_{R}(L,M_{s}),\operatorname{cx}_{R}(L,M_{t})\}$ and $\operatorname{curv}_{R}(L,M_{u})\leqslant\sup\{\operatorname{curv}_{R}(L,M_{s}),\operatorname{curv}_{R}(L,M_{t})\}$.
Proof.
(1) Consider the corresponding long exact sequence of Tor modules. Hence, the first claim follows from 2.11, while the desired inequalities can be obtained from 2.11, [DGS24, Lem. 3.4] and Lemma 2.2 by computing the lengths of the Tor modules.
Similarly, considering the long exact sequences of Ext modules, we obtain (2) and (3). ∎
Lemma 2.13.
An $R$-module $M$ locally has finite projective dimension on punctured spectrum of $R$ if and only if the completion $\widehat{M}$ locally has finite projective dimension on punctured spectrum of $\widehat{R}$.
Proof.
[…]
Next, we recall the notions of Burch (sub)modules and ideals.
Definition 2.14.
  1. (1)
    [DK23, Def. 3.1] A submodule $M$ of an $R$-module $L$ is said to be Burch if
    $\mathfrak{m}(M:_{L}\mathfrak{m})\neq\mathfrak{m}M$, i.e., $\mathfrak{m}(M:_{L}\mathfrak{m})\nsubseteq\mathfrak{m}M$.
    An $R$-module $M$ is called Burch if it is a Burch submodule of some $R$-module.
  2. (2)
    [DKT20, Def. 2.1] A Burch submodule of $R$ is called a Burch ideal of $R$.
[…]
Lemma 2.15.
If $M$ is a Burch $R$-module, then $\widehat{M}$ is a Burch $\widehat{R}-$module.
A key property of a Burch module of positive depth is the following.
Lemma 2.16.
[DK23, Lem. 3.6] Let $M$ be a Burch submodule of an $R$-module $L$ such that $\operatorname{depth}(M)>0$. Then there exists an $M$-regular element $x\in\mathfrak{m}$ such that $k$ is a direct summand of $M/xM$.
Inherent to the definition of Burch submodule is the role of how a module $M$ embeds in another module $L$. We may wonder how important the role of this bigger module is.
Question 2.17.
Let $f:M\to X$ and $g:M\to Y$ be injective $R$-module homomorphisms. If $f(M)$ is a Burch submodule of $X$, then is $g(M)$ a Burch submodule of $Y$?
In order to give a partial answer to 2.17, we prove the following.
Lemma 2.18.
Let $U,V$ and $W$ be $R$-modules such that $U\subseteq V$ and there exists an injective homomorphism $f:V\to W$. Consider an ideal $I$ of $R$. Then, $f\big((U:_{V}I)\big)=\big(f(U):_{f(V)}I\big)$.
Proof.
First, note that $If\big((U:_{V}I)\big)=f\big(I(U:_{V}I)\big)\subseteq f(U)$. Thus, $f\big((U:_{V}I)\big)\subseteq\big(f(U):_{f(V)}I\big)$. For reverse inclusion, let $w\in\big(f(U):_{f(V)}I\big)$. Then, $w=f(v)$ for some $v\in V$, and $f(Iv)=If(v)=Iw\subseteq f(U)$. Hence, since $f$ is injective, $Iv\subseteq U$, i.e., $v\in(U:_{V}I)$. So $w=f(v)\in f\big((U:_{V}I)\big)$. ∎
In a very special case, the next lemma provides an affirmative answer to 2.17, and this will be used extensively in the next section.
Lemma 2.19.
Let $f:M\to X$ and $h:X\to Y$ be injective $R$-module homomorphisms. If $f(M)$ is a Burch submodule of $X$, then $h(f(M))$ is a Burch submodule of $Y$.
Proof.
Put $U:=f(M)$. Then $U$ is a submodule of $X$, and $h(U)$ is a submodule of $h(X)$. If possible, suppose that $h(U)$ is not a Burch submodule of $Y$. Then, $\mathfrak{m}h(U)=\mathfrak{m}\big(h(U):_{Y}\mathfrak{m}\big)$. Hence the inclusions
$\mathfrak{m}h(U)\subseteq\mathfrak{m}\big(h(U):_{h(X)}\mathfrak{m}\big)\subseteq\mathfrak{m}\big(h(U):_{Y}\mathfrak{m}\big)$
imply that $\mathfrak{m}h(U)=\mathfrak{m}\big(h(U):_{h(X)}\mathfrak{m}\big)=\mathfrak{m}\,h\big((U:_{X}\mathfrak{m})\big)$, where the last equality is obtained from Lemma 2.18. Thus, $h(\mathfrak{m}U)=h\big(\mathfrak{m}(U:_{X}\mathfrak{m})\big)$. Since $h$ is injective, it follows that $\mathfrak{m}U=\mathfrak{m}(U:_{X}\mathfrak{m})$, i.e., $U=f(M)$ is not Burch in $X$, which is a contradiction. So $h(f(M))$ is a Burch submodule of $Y$. ∎
Burch modules have maximal (injective) complexity and curvature.
Proposition 2.20.
[Avr96, Thm. 4], [GS24, Thm. 5.1] Let $M$ be a Burch module. Then
  1. (1)
    $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(k)$ and $\operatorname{inj\,cx}_{R}(M)=\operatorname{cx}_{R}(k)$.
  2. (2)
    $\operatorname{curv}_{R}(M)=\operatorname{curv}_{R}(k)$ and $\operatorname{inj\,curv}_{R}(M)=\operatorname{curv}_{R}(k)$.
The next lemma is important for proving results about Burch modules of depth zero.
Lemma 2.21.
[GS24, Lem. 4.8] Let $S=R[X]_{\langle\mathfrak{m},X\rangle}$. Let $M$ be a submodule of an $R$-module $L$. Set $L^{\prime}:=L[X]_{\langle\mathfrak{m},X\rangle}$ and $M^{\prime}:=(M+XL[X])_{\langle\mathfrak{m},X\rangle}$. Then
  1. (1)
    If $M$ is a Burch submodule of $L$, then $M^{\prime}$ is a Burch $S$-submodule of $L^{\prime}$.
  2. (2)
    $M^{\prime}/XM^{\prime}\cong M\oplus L/M$ as $R$-modules.
Remark 2.22.
In the common hypothesis of all the statements in [GS24, Lem. 4.8], it is assumed that $M$ is a Burch submodule of $L$. Note that the Burch property of $M$ is required for the proof of statement (2) in [GS24, Lem. 4.8], and is not needed elsewhere.
The following lemma ensures that the (injective) complexity and curvature of a module are preserved under the base change from $R$ to $R[X]_{\langle\mathfrak{m},X\rangle}$.
Lemma 2.23.
Let $N$ be an $R$-module. Set $S:=R[X]_{\langle\mathfrak{m},X\rangle}$. Then $S$ is a local ring with the same residue field as that of $R$. We may consider $N$ as an $S$-module via the isomorphism $S/XS\cong R$. Moreover, the following hold.
  1. (1)
    $\operatorname{cx}_{S}(N)=\operatorname{cx}_{R}(N)$ and $\operatorname{curv}_{S}(N)=\operatorname{curv}_{R}(N)$.
  2. (2)
    $\operatorname{inj\,cx}_{S}(N)=\operatorname{inj\,cx}_{R}(N)$ and $\operatorname{inj\,curv}_{S}(N)=\operatorname{inj\,curv}_{R}(N)$.
Proof.
Set $N^{\prime}:=N[X]_{\langle\mathfrak{m},X\rangle}$. Note that $X$ is $N^{\prime}$-regular, and $N^{\prime}/XN^{\prime}\cong N$.
(1) The first equality is obtained as follows:
[…]
$=\operatorname{cx}_{R}(N).$
The same equalities hold for curvature as well.
(2) Similarly, as in (1), the desired equalities on injective complexity and curvature can be obtained by using [DGS24, Lem. 3.11.(2)] and Lemma 2.8.(2). ∎
For a submodule $M$ of $L$, we now see how Ext and Tor modules involving $M$ behave if $M$ over $R$ is changed to $M^{\prime}:=(M+XL[X])_{\langle\mathfrak{m},X\rangle}$ over $R[X]_{\langle\mathfrak{m},X\rangle}$.
Lemma 2.24.
Let $M$ be a submodule of an $R$-module $L$, and let $N$ be an $R$-module. Set $S:=R[X]_{\langle\mathfrak{m},X\rangle}$, $L^{\prime}:=L[X]_{\langle\mathfrak{m},X\rangle}$ and $M^{\prime}:=(M+XL[X])_{\langle\mathfrak{m},X\rangle}$. We may consider $N$ as an $S$-module via the isomorphism $S/XS\cong R$. Then we have the following $R$-module isomorphisms.
  1. (1)
    $\operatorname{Tor}_{n}^{S}(M^{\prime},N)\cong\operatorname{Tor}_{n}^{R}(M,N)\oplus\operatorname{Tor}_{n}^{R}(L/M,N)$ for every $n\geqslant 0$.
  2. (2)
    $\operatorname{Ext}^{n}_{S}(M^{\prime},N)\cong\operatorname{Ext}^{n}_{R}(M,N)\oplus\operatorname{Ext}^{n}_{R}(L/M,N)$ for every $n\geqslant 0$.
  3. (3)
    $\operatorname{Ext}^{n+1}_{S}(N,M^{\prime})\cong\operatorname{Ext}^{n}_{R}(N,M)\oplus\operatorname{Ext}^{n}_{R}(N,L/M)$ for every $n\geqslant 0$.
Proof.
Note that $X$ is a regular element on $S$ and $M^{\prime}$. Therefore
[…]
$\cong\operatorname{Tor}_{n}^{R}(M,N)\oplus\operatorname{Tor}_{n}^{R}(L/M,N)\mbox{ for all $n\geqslant 0$}.$
Replacing Tor by Ext, similar isomorphisms as in (2.1) yield (2). Since
[…]
$\cong\operatorname{Ext}^{n}_{R}(N,M)\oplus\operatorname{Ext}^{n}_{R}(N,L/M)\mbox{ for all $n\geqslant 0$},$
the proof of (3) follows. ∎
Some consequences of Lemma 2.24 are the following.
Lemma 2.25.
With the hypotheses as in Lemma 2.24, the following hold.
  1. (1)
    $\operatorname{tcx}_{S}(M^{\prime},N)=\sup\{\operatorname{tcx}_{R}(M,N),\operatorname{tcx}_{R}(L/M,N)\}$.
  2. (2)
    $\operatorname{tcurv}_{S}(M^{\prime},N)=\sup\{\operatorname{tcurv}_{R}(M,N),\operatorname{tcurv}_{R}(L/M,N)\}$.
  3. (3)
    $\operatorname{cx}_{S}(M^{\prime},N)=\sup\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(L/M,N)\}$.
  4. (4)
    $\operatorname{curv}_{S}(M^{\prime},N)=\sup\{\operatorname{curv}_{R}(M,N),\operatorname{curv}_{R}(L/M,N)\}$.
  5. (5)
    $\operatorname{cx}_{S}(N,M^{\prime})=\sup\{\operatorname{cx}_{R}(N,M),\operatorname{cx}_{R}(N,L/M)\}$.
  6. (6)
    $\operatorname{curv}_{S}(N,M^{\prime})=\sup\{\operatorname{curv}_{R}(N,M),\operatorname{curv}_{R}(N,L/M)\}$.
Proof.
Since for any two $R$-modules $U$ and $V$, one has $\mu_{R}(U\oplus V)=\mu_{R}(U)+\mu_{R}(V)$, all the equalities are direct consequences of Lemmas 2.2 and 2.24 simultaneously. ∎
2.26.
Let $\mathcal{A}$ be the category of $R$-modules of Gorenstein dimension zero, and $\mathcal{B}$ be the subcategory of projective $R$-modules. Following the notation as in [Aus-buch, p. 9], $\hat{\mathcal{A}}$ and $\hat{\mathcal{B}}$ denote the category of $R$-modules of finite Gorenstein dimension and of finite projective dimension, respectively. Then, in view of [Aus-buch, Thm. 1.1], for each $U$ in $\hat{\mathcal{A}}$, there is an exact sequence $0\rightarrow U\rightarrow V\rightarrow W\rightarrow 0$ with $V$ in $\hat{\mathcal{B}}$ and $W$ in $\mathcal{A}$.
3. Main results
Here, we provide the main results of this article. We start with the following proposition.
Proposition 3.1.
Let $M$ and $N$ be $R$-modules such that $M$ is Burch. Let $\operatorname{depth}(M)\geqslant 1$, and $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$ for some $h\geqslant 1$. Then
  1. (1)
    $\operatorname{cx}_{R}(N)\leqslant\operatorname{tcx}_{R}(M,N)$ and $\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$. When $R$ is a complete intersection, both inequalities become equalities.
  2. (2)
    Suppose $N$ is also a Burch module. Then the following are equivalent.
    1. (i)
      $R$ is complete intersection ((resp., of codimension $c$)).
    2. (ii)
      $\operatorname{tcx}_{R}(M,N)<\infty$ ((resp., $\operatorname{tcx}_{R}(M,N)=c<\infty$)).
    3. (iii)
      $\operatorname{tcurv}_{R}(M,N)\leqslant 1$.
Proof.
(1) By Lemma 2.16, there exists an $M$-regular element $x\in\mathfrak{m}$ such that $k$ is a direct summand of $M/xM$. Therefore $\operatorname{cx}_{R}(N)=\operatorname{tcx}_{R}(k,N)\leqslant\operatorname{tcx}_{R}(M/xM,N)\leqslant\operatorname{tcx}_{R}(M,N)$, where the last two inequalities are obtained from Lemma 2.10.(1) and [DGS24, Lem. 3.12.(2)] respectively. Similarly, one obtains $\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$.
Suppose that $R$ is complete intersection. Then, by Proposition 2.9.(2), one has $\operatorname{tcx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(N)$. Hence $\operatorname{cx}_{R}(N)=\operatorname{tcx}_{R}(M,N)$. For equality on the curvatures, since $0\leqslant\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$, there is nothing to prove when $\operatorname{tcurv}_{R}(M,N)=0$. So we may assume that $\operatorname{tcurv}_{R}(M,N)>0$, equivalently, $\operatorname{tcx}_{R}(M,N)>0$, cf. Remark 2.6.(3).(d). Note that $\operatorname{tcx}_{R}(M,N)=\operatorname{cx}_{R}(N)<\infty$ (by Proposition 2.7). Hence, by Remark 2.6.(3), $0<\operatorname{tcurv}_{R}(M,N)\leqslant 1$, i.e., $\operatorname{tcurv}_{R}(M,N)=1$. It remains to prove that $\operatorname{curv}_{R}(N)=1$. Since $R$ is complete intersection, $\operatorname{curv}_{R}(N)\leqslant 1$, i.e., $\operatorname{curv}_{R}(N)$ can be 0 or $1$. Since $\operatorname{cx}_{R}(N)=\operatorname{tcx}_{R}(M,N)>0$, it follows that $\operatorname{curv}_{R}(N)=1$. Thus, $\operatorname{curv}_{R}(N)=\operatorname{tcurv}_{R}(M,N)$.
(2) First we note that if $R$ is complete intersection, in view of Proposition 2.7.(2),
(3.1) $\operatorname{codim}(R)=\operatorname{cx}_{R}(k)=\operatorname{cx}_{R}(N)=\operatorname{tcx}_{R}(M,N),$
where the last two equalities are obtained from Proposition 2.20.(1) and (1) respectively. So, the implications (i) $\Rightarrow$ (ii) and (i) $\Rightarrow$ (iii) follow from Proposition 2.9.(2), while the implication (ii) $\Rightarrow$ (iii) is trivial (Remark 2.6.(3).(b)). Assuming (iii), by Proposition 2.20.(2) and (1), $\operatorname{curv}_{R}(k)=\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)\leqslant 1$, and hence $R$ is complete intersection by Proposition 2.7.(2). Thus, one also sees that if $\operatorname{tcx}_{R}(M,N)=c<\infty$, then $R$ is complete intersection of codimension $c$, see (3.1). ∎
For Burch modules of arbitrary depth, we prove the following result on Tor complexity and curvature.
Theorem 3.2.
Let $L,M$ and $N$ be $R$-modules such that $M$ is a Burch submodule of $L$. Let $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(L,N)=0$ for some $h\geqslant 1$. Then
  1. (1)
    $\operatorname{cx}_{R}(N)\leqslant\max\{\operatorname{tcx}_{R}(M,N),\operatorname{tcx}_{R}(L/M,N)\}$.
  2. (2)
    $\operatorname{curv}_{R}(N)\leqslant\max\{\operatorname{tcurv}_{R}(M,N),\operatorname{tcurv}_{R}(L/M,N)\}$.
  3. (3)
    Both inequalities (1) and (2) become equalities if $R$ is a complete intersection.
  4. (4)
    Suppose $N$ is also a Burch module. Then the following are equivalent.
    1. (i)
      $R$ is complete intersection.
    2. (ii)
      $\max\{\operatorname{tcx}_{R}(M,N),\operatorname{tcx}_{R}(L/M,N)\}<\infty$.
    3. (iii)
      $\max\{\operatorname{tcurv}_{R}(M,N),\operatorname{tcurv}_{R}(L/M,N)\}\leqslant 1$.
    Under these conditions, $\operatorname{codim}(R)=\max\{\operatorname{tcx}_{R}(M,N),\operatorname{tcx}_{R}(L/M,N)\}$.
Proof.
(1) When $\operatorname{depth}(M)\geqslant 1$, the inequality follows from Proposition 3.1.(1). So, we may assume that $\operatorname{depth}(M)=0$. Consider $S:=R[X]_{\langle\mathfrak{m},X\rangle}$, $L^{\prime}:=L[X]_{\langle\mathfrak{m},X\rangle}$ and $M^{\prime}:=(M+XL[X])_{\langle\mathfrak{m},X\rangle}$. We may view $N$ as an $S$-module via the isomorphism $S/XS\cong R$. Since $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$ and $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(L,N)=0$, it follows from Lemma 2.12.(1) that $\mathfrak{m}^{2h}\operatorname{Tor}^{R}_{\gg 0}(L/M,N)=0$. Hence, in view of Lemma 2.24.(1), $\mathfrak{n}^{2h}\operatorname{Tor}^{S}_{\gg 0}(M^{\prime},N)=0$, where $\mathfrak{n}$ is the maximal ideal of $S$. Note that $M^{\prime}$ is a Burch $S$-submodule of $L^{\prime}$ by Lemma 2.21.(1). Therefore, since $\operatorname{depth}_{S}(M^{\prime})\geqslant 1$, by Proposition 3.1.(1), $\operatorname{cx}_{S}(N)\leqslant\operatorname{tcx}_{S}(M^{\prime},N)$. Due to Lemma 2.23.(1), one has $\operatorname{cx}_{S}(N)=\operatorname{cx}_{R}(N)$. Hence, by Lemma 2.25.(1), $\operatorname{cx}_{R}(N)\leqslant\operatorname{tcx}_{S}(M^{\prime},N)=\max\{\operatorname{tcx}_{R}(M,N),\operatorname{tcx}_{R}(L/M,N)\}$.
(2) Replacing $\operatorname{cx}$ by $\operatorname{curv}$, the proof follows along the same lines as (1).
(3) Assume that $R$ is a complete intersection. Then, by Proposition 2.9.(2), $\operatorname{tcx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(N)$. Similarly, $\operatorname{tcx}_{R}(L/M,N)\leqslant\operatorname{cx}_{R}(N)$. Thus,
$\max\{\operatorname{tcx}_{R}(M,N),\operatorname{tcx}_{R}(L/M,N)\}\leqslant\operatorname{cx}_{R}(N).$
Hence, by (1), the equality follows. Analogous argument shows the equality for curvature.
(4) In view of Proposition 2.20, $\operatorname{cx}_{R}(N)=\operatorname{cx}_{R}(k)$ and $\operatorname{curv}_{R}(N)=\operatorname{curv}_{R}(k)$. So the desired equivalences follow from (1), (2), Proposition 2.7.(2) and Proposition 2.9.(2). Hence the last part can be obtained from (3) and Proposition 2.7.(2). ∎
The following is a counterpart of Proposition 3.1 for Ext complexity and curvature.
Proposition 3.3.
Let $M$ and $N$ be $R$-modules such that $M$ is Burch. Let $\operatorname{depth}(M)\geqslant 1$, and $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(M,N)=0$ for some $h\geqslant 1$. Then
  1. (1)
    $\operatorname{inj\,cx}_{R}(N)\leqslant\operatorname{cx}_{R}(M,N)$ and $\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$. When $R$ is a complete intersection, both inequalities become equalities.
  2. (2)
    Assume $N$ is also a Burch module. Then the following conditions are equivalent.
    1. (i)
      $R$ is complete intersection ((resp., of codimension $c$)).
    2. (ii)
      $\operatorname{cx}_{R}(M,N)<\infty$ ((resp., $\operatorname{cx}_{R}(M,N)=c<\infty$)).
    3. (iii)
      $\operatorname{curv}_{R}(M,N)\leqslant 1$.
Proof.
(1) By Lemma 2.16, there exists an $M$-regular element $x$ of $R$ such that $k$ is a direct summand of $M/xM$. Therefore $\operatorname{inj\,cx}_{R}(N)=\operatorname{cx}_{R}(k,N)\leqslant\operatorname{cx}_{R}(M/xM,N)\leqslant\operatorname{cx}_{R}(M,N)$, where the last two inequalities are obtained from Lemma 2.10.(3) and [DGS24, Lem. 3.13.(2)] respectively. Similarly, one can obtain $\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$.
When $R$ is complete intersection, due to Proposition 2.9.(1), $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(N)=\operatorname{inj\,cx}_{R}(N)$. Hence $\operatorname{inj\,cx}_{R}(N)=\operatorname{cx}_{R}(M,N)$. For equality on the curvatures, since $0\leqslant\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$, there is nothing to prove when $\operatorname{curv}_{R}(M,N)=0$. So we may assume that $\operatorname{curv}_{R}(M,N)>0$, equivalently, $\operatorname{cx}_{R}(M,N)>0$. Therefore $\operatorname{cx}_{R}(M,N)=\operatorname{inj\,cx}_{R}(N)<\infty$ by Proposition 2.7(1).(i) and 2.7.(2). Hence, due to Remark 2.6.(3), $\operatorname{curv}_{R}(M,N)\leqslant 1$. Thus $\operatorname{curv}_{R}(M,N)=1$. It remains to prove that $\operatorname{inj\,curv}_{R}(N)=1$. Since $R$ is complete intersection, $\operatorname{inj\,curv}_{R}(N)=\operatorname{curv}_{R}(N)\leqslant 1$, i.e., $\operatorname{inj\,curv}_{R}(N)$ can be 0 or $1$. Since $\operatorname{inj\,cx}_{R}(N)=\operatorname{cx}_{R}(M,N)>0$, it follows that $\operatorname{inj\,curv}_{R}(N)=1$.
(2) When $R$ is complete intersection, in view of Proposition 2.7.(2),
(3.2) $\operatorname{codim}(R)=\operatorname{cx}_{R}(k)=\operatorname{inj\,cx}_{R}(N)=\operatorname{cx}_{R}(M,N),$
where the last two equalities hold due to Proposition 2.20.(1) and (1) respectively. So, the implications (i) $\Rightarrow$ (ii) and (i) $\Rightarrow$ (iii) follow from Proposition 2.9.(1), whereas (ii) $\Rightarrow$ (iii) is trivial (Remark 2.6.(3).(a)). Assuming (iii), by Proposition 2.20.(2) and (1) $\operatorname{curv}_{R}(k)=\operatorname{curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)\leqslant 1$. Hence $R$ is complete intersection by Proposition 2.7.(2). Thus, one also sees that if $\operatorname{cx}_{R}(M,N)=c<\infty$, then $R$ is complete intersection of codimension $c$, see (3.2). ∎
The following, which can be thought of as an Ext-counterpart to Theorem 3.2, shows that when $M$ is a Burch submodule of $L$ of depth zero, in addition to $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(M,N)=0$, if $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(L,N)=0$ for some $h\geqslant 1$, then the following inequalities hold.
Theorem 3.4.
Let $L,M$ and $N$ be $R$-modules such that $M$ is a Burch submodule of $L$. Suppose that $\mathfrak{m}^{h}\operatorname{Ext}^{\gg 0}_{R}(M,N)=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(L,N)=0$ for some $h\geqslant 1$. Then
  1. (1)
    $\operatorname{inj\,cx}_{R}(N)\leqslant\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(L/M,N)\}$.
  2. (2)
    $\operatorname{inj\,curv}_{R}(N)\leqslant\max\{\operatorname{curv}_{R}(M,N),\operatorname{curv}_{R}(L/M,N)\}$.
  3. (3)
    Both inequalities (1) and (2) become equalities if $R$ is a complete intersection.
  4. (4)
    Assume $N$ is also a Burch module. Then the following conditions are equivalent.
    1. (i)
      $R$ is complete intersection.
    2. (ii)
      $\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(L/M,N)\}<\infty$.
    3. (iii)
      $\max\{\operatorname{curv}_{R}(M,N),\operatorname{curv}_{R}(L/M,N)\}\leqslant 1$.
    Under these conditions, $\operatorname{codim}(R)=\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(L/M,N)\}$.
Proof.
(1) When $\operatorname{depth}(M)\geqslant 1$, the inequality follows from Proposition 3.3.(1). So assume that $\operatorname{depth}(M)=0$. Set $S$, $L^{\prime}$ and $M^{\prime}$ as defined in the proof of Theorem 3.2.(1), and consider $N$ as an $S$-module via the isomorphism $S/XS\cong R$. Since $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ and $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(L,N)=0$, we obtain $\mathfrak{m}^{2h}\operatorname{Ext}_{R}^{\gg 0}(L/M,N)=0$ by Lemma 2.12.(2). Hence, in view of Lemma 2.24.(2), $\mathfrak{n}^{2h}\operatorname{Ext}_{S}^{\gg 0}(M^{\prime},N)=0$, where $\mathfrak{n}$ is the maximal ideal of $S$. Note that $M^{\prime}$ is a Burch $S$-submodule of $L^{\prime}$ by Lemma 2.21.(1). Therefore, since $\operatorname{depth}_{S}(M^{\prime})\geqslant 1$, by Proposition 3.3.(1), $\operatorname{inj\,cx}_{S}(N)\leqslant\operatorname{cx}_{S}(M^{\prime},N)$. Due to Lemma 2.23.(2), $\operatorname{inj\,cx}_{S}(N)=\operatorname{inj\,cx}_{R}(N)$. Hence, by Lemma 2.25.(3), $\operatorname{inj\,cx}_{R}(N)\leqslant\operatorname{cx}_{S}(M^{\prime},N)=\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(L/M,N)\}$.
(2) Replacing $\operatorname{cx}$ by $\operatorname{curv}$, the proof follows along the same lines as (1).
(3) Assume that $R$ is complete intersection. Then, by Proposition 2.9(1), we obtain $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(N)=\operatorname{inj\,cx}_{R}(N)$. Similarly, $\operatorname{cx}_{R}(L/M,N)\leqslant\operatorname{inj\,cx}_{R}(N)$. Thus $\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(L/M,N)\}\leqslant\operatorname{inj\,cx}_{R}(N)$. Hence, by (1), the equality follows. Analogous argument shows the equality for curvature.
(4) In view of Proposition 2.20, we have $\operatorname{inj\,cx}_{R}(N)=\operatorname{cx}_{R}(k)$ and $\operatorname{inj\,curv}_{R}(N)=\operatorname{curv}_{R}(k)$. So, the desired equivalences follow from (1), (2), Proposition 2.7.(2) and Proposition 2.9.(1). Hence, the last part can be obtained from (3) and Proposition 2.7.(2). ∎
The following result is an analogue of Proposition 3.3, where $N$ is assumed to be Burch in place of $M$.
Proposition 3.5.
Let $M$ and $N$ be $R$-modules such that $N$ is Burch. Let $\operatorname{depth}(N)\geqslant 1$, and $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ for some $h\geqslant 1$. Then
  1. (1)
    $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,N)$ and $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$. When $R$ is complete intersection, both the inequalities become equalities.
  2. (2)
    Suppose $M$ is also a Burch module. Then the following are equivalent.
    1. (i)
      $R$ is complete intersection ((resp., of codimension $c$)).
    2. (ii)
      $\operatorname{cx}_{R}(M,N)<\infty$ ((resp., $\operatorname{cx}_{R}(M,N)=c<\infty$)).
    3. (iii)
      $\operatorname{curv}_{R}(M,N)\leqslant 1$.
Proof.
(1) By Lemma 2.16, there exists an $N$-regular element $x$ of $R$ such that $k$ is a direct summand of $N/xN$. Therefore $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(M,k)\leqslant\operatorname{cx}_{R}(M,N/xN)\leqslant\operatorname{cx}_{R}(M,N)$, where the last two inequalities are obtained by using Lemma 2.10.(5) and [DGS24, Lem. 3.13(1)] respectively. The inequality for curvature can be obtained using similar arguments.
When $R$ is complete intersection, due to Proposition 2.9.(1), $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(M)$. Hence $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(M,N)$. For equality on the curvatures, since $0\leqslant\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$, there is nothing to prove when $\operatorname{curv}_{R}(M,N)=0$. So we may assume that $\operatorname{curv}_{R}(M,N)>0$, equivalently, $\operatorname{cx}_{R}(M,N)>0$. Therefore $\operatorname{cx}_{R}(M,N)=\operatorname{cx}_{R}(M)<\infty$ by Proposition 2.7(1).(i) and 2.7.(2). Hence, due to Remark 2.6.(3), $\operatorname{curv}_{R}(M,N)\leqslant 1$. Thus $\operatorname{curv}_{R}(M,N)=1$. It remains to prove that $\operatorname{curv}_{R}(M)=1$. Since $R$ is complete intersection, $\operatorname{curv}_{R}(M)\leqslant 1$, i.e., $\operatorname{curv}_{R}(M)$ can be 0 or $1$. Since $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(M,N)>0$, it follows that $\operatorname{curv}_{R}(M)=1$.
(2) When $R$ is complete intersection, in view of Proposition 2.7.(2),
(3.3) $\operatorname{codim}(R)=\operatorname{cx}_{R}(k)=\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(M,N),$
where the last two equalities hold due to Proposition 2.20.(1) and (1) respectively. So, the implications (i) $\Rightarrow$ (ii) and (i) $\Rightarrow$ (iii) follow from Proposition 2.9.(1), whereas (ii) $\Rightarrow$ (iii) is trivial (Remark 2.6.(3).(a)). Now, assuming (iii), by Proposition 2.20.(2) and (1) $\operatorname{curv}_{R}(k)=\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)\leqslant 1$. Hence $R$ is complete intersection by Proposition 2.7.(2). Thus, one also sees that if $\operatorname{cx}_{R}(M,N)=c<\infty$, then $R$ is complete intersection of codimension $c$, see (3.3). ∎
Under the assumptions that $N$ is a Burch submodule of $L$ and $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,L)=0$ for some $h\geqslant 1$, the following result is analogous to Theorem 3.4.
Theorem 3.6.
Let $L,M$ and $N$ be $R$-modules such that $N$ is a Burch submodule of $L$. Let $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,L)=0$ for some $h\geqslant 1$. Then
  1. (1)
    $\operatorname{cx}_{R}(M)\leqslant\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(M,L/N)\}$.
  2. (2)
    $\operatorname{curv}_{R}(M)\leqslant\max\{\operatorname{curv}_{R}(M,N),\operatorname{curv}_{R}(M,L/N)\}$.
  3. (3)
    Both inequalities (1) and (2) become equalities if $R$ is a complete intersection.
  4. (4)
    Suppose $M$ is also a Burch module. Then the following are equivalent.
    1. (i)
      $R$ is complete intersection.
    2. (ii)
      $\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(M,L/N)\}<\infty$.
    3. (iii)
      $\max\{\operatorname{curv}_{R}(M,N),\operatorname{curv}_{R}(M,L/N)\}\leqslant 1$.
    Under these conditions, $\operatorname{codim}(R)=\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(M,L/N)\}$.
Proof.
(1) When $\operatorname{depth}(N)\geqslant 1$, the inequality follows from Proposition 3.5.(1). Therefore, assume that $\operatorname{depth}(N)=0$. Consider $S:=R[X]_{\langle\mathfrak{m},X\rangle}$, $L^{\prime}:=L[X]_{\langle\mathfrak{m},X\rangle}$ and $N^{\prime}:=(N+XL[X])_{\langle\mathfrak{m},X\rangle}$. We may view $M$ as $S$-module via the isomorphism $S/XS\cong R$. Since $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ and $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,L)=0$, it follows from Lemma 2.12.(3) that $\mathfrak{m}^{2h}\operatorname{Ext}_{R}^{\gg 0}(M,L/N)=0$. Therefore, by Lemma 2.24(3), we get $\mathfrak{n}^{2h}\operatorname{Ext}_{R}^{\gg 0}(M,N^{\prime})=0$, where $\mathfrak{n}$ is the maximal ideal of $S$. Note that $N^{\prime}$ is a Burch $S$-submodule of $L^{\prime}$ by Lemma 2.21.(1). Therefore, since $\operatorname{depth}_{S}(N^{\prime})\geqslant 1$, by Proposition 3.5.(1), $\operatorname{cx}_{S}(M)\leqslant\operatorname{cx}_{S}(M,N^{\prime})$. Due to Lemma 2.23.(1), $\operatorname{cx}_{S}(M)=\operatorname{cx}_{R}(M)$. Hence, by Lemma 2.25.(5), $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{S}(M,N^{\prime})=\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(M,L/N)\}$.
(2) Replacing $\operatorname{cx}$ by $\operatorname{curv}$, the proof follows along the same lines as (1).
(3) Now, let $R$ be complete intersection. From Proposition 2.9(1), $\operatorname{cx}_{R}(M,N)\leqslant\min\{\operatorname{cx}_{R}(M),\operatorname{cx}_{R}(N)\}\leqslant\operatorname{cx}_{R}(M)$. Similarly, $\operatorname{cx}_{R}(M,L/N)\leqslant\operatorname{cx}_{R}(M)$. Thus
$\max\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(M,L/N)\}\leqslant\operatorname{cx}_{R}(M).$
Hence, by (1), the equality follows. The analogous argument shows the equality for curvature.
(4) In view of Proposition 2.20, we have $\operatorname{cx}_{R}(M)=\operatorname{cx}_{R}(k)$ and $\operatorname{curv}_{R}(M)=\operatorname{curv}_{R}(k)$. So the desired equivalences follow from (1), (2), Proposition 2.7.(2) and Proposition 2.9.(1). Hence the last part can be obtained using (3) and Proposition 2.7.(2). ∎
The next three propositions show that, under certain conditions, when one of the modules in the pair $(M,N)$ is Burch, the (injective) complexity (resp., curvature) of the other module is bounded above by the Ext or Tor complexity (resp., curvature) of the pair.
Proposition 3.7.
Let $L,M$ and $N$ be $R$-modules such that $M$ is a Burch submodule of $L$. Suppose that $\mathfrak{m}^{h}\operatorname{Tor}_{\gg 0}^{R}(L,N)=0$ and $\mathfrak{m}^{h}\operatorname{Tor}_{\gg 0}^{R}(M,N)=0$ for some $h>0$.
  1. (1)
    If $\operatorname{tcx}_{R}(L,N)\leqslant\operatorname{tcx}_{R}(M,N)$, then $\operatorname{cx}_{R}(N)\leqslant\operatorname{tcx}_{R}(M,N)$.
  2. (2)
    If $\operatorname{tcurv}_{R}(L,N)\leqslant\operatorname{tcurv}_{R}(M,N)$, then $\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$.
Proof.
In view of Lemma 2.12.(1), the short exact sequence $0\to M\to L\to L/M\to 0$ yields $\operatorname{tcx}_{R}(L/M,N)\leqslant\sup\{\operatorname{tcx}_{R}(M,N),\operatorname{tcx}_{R}(L,N)\}=\operatorname{tcx}_{R}(M,N)$ and $\operatorname{tcurv}_{R}(L/M,N)\leqslant\sup\{\operatorname{tcurv}_{R}(M,N),\operatorname{tcurv}_{R}(L,N)\}=\operatorname{tcurv}_{R}(M,N)$. Hence, we are done by Theorem 3.2.(1) and Theorem 3.2.(2). ∎
Proposition 3.8.
Let $L,M$ and $N$ be $R$-modules such that $M$ is a Burch submodule of $L$. Suppose that $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(L,N)=0$ and $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ for some $h>0$.
  1. (1)
    If $\operatorname{cx}_{R}(L,N)\leqslant\operatorname{cx}_{R}(M,N)$, then $\operatorname{inj\,cx}_{R}(N)\leqslant\operatorname{cx}_{R}(M,N)$.
  2. (2)
    If $\operatorname{curv}_{R}(L,N)\leqslant\operatorname{curv}_{R}(M,N)$, then $\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$.
Proof.
The proof is similar to Proposition 3.7 using Lemma 2.12.(2) and Theorem 3.4 in places of Lemma 2.12.(1) and Theorem 3.2 respectively. ∎
Proposition 3.9.
Let $L,M$ and $N$ be $R$-modules such that $N$ is a Burch submodule of $L$. Suppose that $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,L)=0$ and $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ for some $h>0$.
  1. (1)
    If $\operatorname{cx}_{R}(M,L)\leqslant\operatorname{cx}_{R}(M,N)$, then $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,N)$.
  2. (2)
    If $\operatorname{curv}_{R}(M,L)\leqslant\operatorname{curv}_{R}(M,N)$, then $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$.
Proof.
In view of Lemma 2.12.(3), the short exact sequence $0\to N\to L\to L/N\to 0$ yields $\operatorname{cx}_{R}(M,L/N)\leqslant\sup\{\operatorname{cx}_{R}(M,N),\operatorname{cx}_{R}(M,L)\}=\operatorname{cx}_{R}(M,N)$ and $\operatorname{curv}_{R}(M,L/N)\leqslant\sup\{\operatorname{curv}_{R}(M,N),\operatorname{curv}_{R}(M,L)\}=\operatorname{curv}_{R}(M,N)$. Hence, we are done by Theorem 3.6.(1) and Theorem 3.6.(2). ∎
The next result shows that if the base ring is CM, $N$ is Burch, and some power of the maximal ideal annihilates $\operatorname{Ext}_{R}^{\gg 0}(M,N)$, then the complexity (resp., curvature) of $M$ is at most that of the pair $(M,N)$.
Theorem 3.10.
Assume that $R$ is CM. Let $M$ and $N$ be $R$-modules such that $N$ is Burch, and $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ for some $h>0$. Then $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,N)$ and $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$.
Proof.
All hypotheses and conclusions are preserved under passing to completion (for the Burch property, see Lemma 2.15). Hence, we may assume that $R$ is complete. Let $N$ be a Burch submodule of an $R$-module $X$. Since $R$ is complete and CM, there exist a finitely generated $R$-module $L$ of finite injective dimension and an injective homomorphism $h:X\to L$ (see [lw, 11.8 and 11.17]). By Lemma 2.19, $h(N)$ is a Burch submodule of $L$. Note that $\operatorname{Ext}^{\gg 0}_{R}(M,L)=0$ (i.e., $\operatorname{cx}_{R}(M,L)=0$), and $h(N)\cong N$. Therefore, by Proposition 3.9, $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,h(N))=\operatorname{cx}_{R}(M,N)$, and similarly $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$. ∎
Before we proceed to the next result, we require the following preparatory lemma.
Lemma 3.11.
Suppose that $R$ is CM of dimension $d$, and it admits a canonical module. Let $M$ and $N$ be $R$-modules such that $N$ locally has finite injective dimension on the punctured spectrum. Then, there exist MCM $R$-modules $X$ and $Y$ such that $Y$ locally has finite injective dimension on punctured spectrum, and $\operatorname{Ext}^{n+d}_{R}(M,N)\cong\operatorname{Ext}_{R}^{n}(X,Y)$ for all $n\geqslant 1$.
Proof.
Set $X:=\operatorname{\Omega}^{d}_{R}(M)$. Then $X$ is MCM, and $\operatorname{Ext}^{n+d}_{R}(M,N)\cong\operatorname{Ext}_{R}^{n}(X,N)$ for all $n\geqslant 1$. Consider an MCM approximation (cf. [lw, 11.8 and 11.17]) of $N$, i.e., a short exact sequence $0\to L\to Y\to N\to 0$, where $L$ has finite injective dimension, and $Y$ is MCM. Hence, the desired isomorphisms follow because $\operatorname{Ext}_{R}^{n}(X,L)=0$ for all $n\geqslant 1$. Note that $Y$ locally has finite injective dimension on punctured spectrum as both $L$ and $N$ have the same property. ∎
When the completion $\widehat{N}$ locally has finite injective dimension on the punctured spectrum of $\widehat{R}$, we establish a result similar to Theorem 3.10.
Proposition 3.12.
Suppose that $R$ is CM. Let $M$ and $N$ be $R$-modules such that $N$ is Burch. If $\widehat{N}$ locally has finite injective dimension on the punctured spectrum of $\widehat{R}$, then $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,N)$ and $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$.
Proof.
Set $d:=\dim(R)$. As in the proof of Theorem 3.10, passing to completion, we may assume that $R$ is complete, and $N$ locally has finite injective dimension on the punctured spectrum. Hence, since $R$ is CM, it admits a canonical module, say $\omega$. Set $(-)^{\dagger}:=\operatorname{Hom}_{R}(-,\omega)$. By Lemma 3.11, there exist MCM $R$-modules $X$ and $Y$ such that $Y$ locally has finite injective dimension on the punctured spectrum, and $\operatorname{Ext}^{n+d}_{R}(M,N)\cong\operatorname{Ext}_{R}^{n}(X,Y)\cong\operatorname{Ext}_{R}^{n}(Y^{\dagger},X^{\dagger})$ for all $n\geqslant 1$, where the last isomorphism is by duality of dualizing module, see, e.g., [cluster, 3.14 and 3.15]. Since $Y$ is MCM and locally has finite injective dimension on the punctured spectrum, for every non-maximal prime ideal $\mathfrak{p}$ of $R$, the localization $Y_{\mathfrak{p}}$ is a finite direct sum of copies of $\omega_{\mathfrak{p}}$. Thus, $Y^{\dagger}$ is locally free on the punctured spectrum of $R$. Therefore, by [DGS24, Lem. 3.6], there exists $h>0$ such that $0=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{n}(Y^{\dagger},X^{\dagger})=\mathfrak{m}^{h}\operatorname{Ext}_{R}^{n+d}(M,N)$ for all $n\gg 0$. Hence, the proposition follows from Theorem 3.10. ∎
As a consequence of the above proposition, we obtain the following.
Theorem 3.13.
Let $R$ be CM such that $\widehat{R}$ is locally Gorenstein on the punctured spectrum. Let $M$ and $N$ be $R$-modules such that $N$ is Burch. Assume that $N$ locally has finite projective dimension on the punctured spectrum. Then $\operatorname{cx}_{R}(M)\leqslant\operatorname{cx}_{R}(M,N)$ and $\operatorname{curv}_{R}(M)\leqslant\operatorname{curv}_{R}(M,N)$.
Proof.
In view of Lemma 2.13, $\widehat{N}$ locally has finite projective dimension on the punctured spectrum of $\widehat{R}$. Hence, since $\widehat{R}$ is locally Gorenstein on the punctured spectrum, $\widehat{N}$ locally has finite injective dimension on the punctured spectrum of $\widehat{R}$. Therefore the proof follows from Proposition 3.12. ∎
In general, it may happen that $R$ is CM with isolated singularity, but $\widehat{R}$ is locally not even Gorenstein on the punctured spectrum, see [heit, Main Theorem]. In the following remark, we note a class of local rings whose completions are locally Gorenstein on the punctured spectrum.
Remark 3.14.
  1. (1)
    If $R$ is Gorenstein, then $\widehat{R}$ is Gorenstein, hence $\widehat{R}$ is locally Gorenstein on the punctured spectrum. If $R$ is Artinian, then trivially $\widehat{R}$ is locally Gorenstein on the punctured spectrum.
  2. (2)
    Suppose that $R$ is CM of dimension $1$. Then it is well known that $\widehat{R}$ is locally Gorenstein on the punctured spectrum if and only if $R$ admits a canonical ideal. Indeed, in view of [agl, Prop. 2.7], $R$ admits a canonical ideal if and only if the total ring of fractions Q(R^)Q($\widehat${R}) is Gorenstein if and only if $\widehat{R}$ is locally Gorenstein at all minimal primes of $\widehat{R}$, where the last equivalence follows from [stacks-project, Lem. 10.25.4] and the fact that the direct product of finitely many rings is Gorenstein if and only if all the components are Gorenstein.
  3. (3)
    If $R$ is CM that admits a canonical module, then $\widehat{R}$ is locally Gorenstein on the punctured spectrum if and only if $R$ is locally Gorenstein on the punctured spectrum; this can be obtained by applying Lemma 2.13 to the canonical module.
When $M$ is Burch submodule of an $R$-module $X$ which either has finite $G$-dimension or is torsionless, the following is an analogue of Theorem 3.10.
Theorem 3.15.
Let $M$ and $N$ be $R$-modules such that $M$ is a Burch submodule of some $R$-module $X$. Assume that $\operatorname{G-dim}_{R}(X)<\infty$, or $X$ is torsionless. Then, the following hold.
  1. (1)
    If $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$ for some $h>0$, then $\operatorname{cx}_{R}(N)\leqslant\operatorname{tcx}_{R}(M,N)$ and $\operatorname{curv}_{R}(N)\leqslant\operatorname{tcurv}_{R}(M,N)$.
  2. (2)
    If $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ for some $h>0$, then $\operatorname{inj\,cx}_{R}(N)\leqslant\operatorname{cx}_{R}(M,N)$ and $\operatorname{inj\,curv}_{R}(N)\leqslant\operatorname{curv}_{R}(M,N)$.
Proof.
First, assume that $\operatorname{G-dim}_{R}(X)<\infty$. Then, in view of 2.26, there exists an $R$-module $L$ of finite projective dimension and an injective homomorphism $h:X\to L$. By Lemma 2.19, $h(M)$ is a Burch submodule of $L$. Note that $\operatorname{Tor}^{R}_{\gg 0}(L,N)=0=\operatorname{Ext}^{\gg 0}_{R}(L,N)$, and $h(M)\cong M$. Therefore, we are done by Propositions 3.7 and 3.8.
For the second case, assume that $X$ is torsionless. Then, there exist a free $R$-module $L$ and an injective homomorphism $h:X\to L$. Hence, the proof is similar as above. ∎
Remark 3.16.
  1. (1)
    The hypotheses $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(M,N)=0$ and/or $\mathfrak{m}^{h}\operatorname{Tor}^{R}_{\gg 0}(L,N)=0$ for some $h\geqslant 1$ in Propositions 3.1, 3.7 and Theorems 3.2, 3.15.(1) hold if $N$ locally has finite projective dimension on the punctured spectrum; see [DGS24, Lem. 3.6.(2)].
  2. (2)
    The hypotheses $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,N)=0$ and/or $\mathfrak{m}^{h}\operatorname{Ext}_{R}^{\gg 0}(M,L)=0$ for some $h\geqslant 1$ in Propositions 3.3, 3.5, 3.9 and Theorems 3.6, 3.10, 3.15.(2) hold if $M$ locally has finite projective dimension on the punctured spectrum; see [DGS24, Lem. 3.6.(1)].
  3. (3)
    Thus, when $R$ has an isolated singularity, the annihilation conditions on Tor/Ext modules in Propositions 3.1, 3.3, 3.5, 3.7, 3.8, 3.9 and Theorems 3.2, 3.4, 3.6, 3.10, 3.15 are immediate.
Although different in spirit from the rest of the results in this article, the following proposition demonstrates the usefulness of Lemma 2.19 by improving an injective dimension test result from [GS24, Cor. 5.4].
Proposition 3.17.
Let $M$ and $N$ be $R$-modules such that $M$ is a Burch submodule of some $R$-module $X$. Suppose $n\geqslant\max\{1,\operatorname{depth}(N)\}$ is an integer such that $\operatorname{Ext}^{i}_{R}(M,N)=0$ for i=n1,n,n+1i=n-1,n,n+1, and at least one of the following holds.
  1. (1)
    $\operatorname{G-dim}_{R}(X)<\infty$, and $n>\operatorname{depth}(R)$.
  2. (2)
    $X$ is torsionless.
Then, $\operatorname{id}_{R}(N)<\infty$.
Proof.
(1) In view of 2.26, there exist an $R$-module $L$ of finite projective dimension and an injective homomorphism $h:X\to L$. By Lemma 2.19, $h(M)$ is a Burch submodule of $L$. Since $n>\operatorname{depth}(R)\geqslant\operatorname{pd}_{R}(L)$, it follows that $\operatorname{Ext}_{R}^{i}(L,N)=0$ for $i=n,n+1$. Since $h(M)\cong M$, by [GS24, Thm. 5.3], $\operatorname{id}_{R}(N)<\infty$.
(2) Since $X$ is torsionless, there exist a free $R$-module $L$ and an injective homomorphism $h:X\to L$. Note that $\operatorname{Ext}_{R}^{\geqslant 1}(L,N)=0$. Hence, as in (1), $\operatorname{id}_{R}(N)<\infty$. ∎
We now provide the proofs of Corollaries 1.5 and 1.6.
Proof of Corollary 1.5.
Note that every $\mathfrak{m}$-primary ideal is locally free on the punctured spectrum. Since $R$ is torsionless, $I$ is $\mathfrak{m}$-primary and Burch, by Theorem 1.4, $\operatorname{cx}_{R}(J)\leqslant\operatorname{tcx}_{R}(I,J)$ and $\operatorname{inj\,cx}_{R}(J)\leqslant\operatorname{cx}_{R}(I,J)$. Therefore, since $J$ is also Burch, in view of Proposition 2.20.(1),
(3.4) $\operatorname{cx}_{R}(k)\leqslant\operatorname{tcx}_{R}(I,J)\;\mbox{ and }\;\operatorname{cx}_{R}(k)\leqslant\operatorname{cx}_{R}(I,J).$
Thus, if $\operatorname{cx}_{R}(k)=\infty$, then clearly $\operatorname{cx}_{R}(k)=\operatorname{tcx}_{R}(I,J)=\operatorname{cx}_{R}(I,J)$. So we may assume that $\operatorname{cx}_{R}(k)$ is finite. Then, by Proposition 2.7.(2), $R$ is complete intersection. Hence, by Propositions 2.9 and 2.7.(1).(i), $\operatorname{tcx}_{R}(I,J)\leqslant\operatorname{cx}_{R}(k)$ and $\operatorname{cx}_{R}(I,J)\leqslant\operatorname{cx}_{R}(k)$. Thus $\operatorname{tcx}_{R}(I,J)=\operatorname{cx}_{R}(I,J)=\operatorname{cx}_{R}(k)$. In view of Proposition 2.7.(2), the equivalence of (1), (2) and (3) follows from these equalities. Arguing similarly as in (3.4), Theorem 1.4 and Proposition 2.20.(2) yield $\operatorname{curv}_{R}(k)\leqslant\operatorname{tcurv}_{R}(I,J)$ and $\operatorname{curv}_{R}(k)\leqslant\operatorname{curv}_{R}(I,J)$. Hence, the implications $(4)\implies(1)$ and $(5)\implies(1)$ follow from Proposition 2.7.(2). The reverse implications are obvious from Proposition 2.9. ∎
Proof of Corollary 1.6.
In view of Proposition 2.20 and Theorem 3.13, $\operatorname{cx}_{R}(k)=\operatorname{cx}_{R}(I)\leqslant\operatorname{cx}_{R}(I,J)$ and $\operatorname{curv}_{R}(k)=\operatorname{curv}_{R}(I)\leqslant\operatorname{curv}_{R}(I,J)$. Thus, if $\operatorname{cx}_{R}(k)=\infty$, then $\operatorname{cx}_{R}(k)=\operatorname{cx}_{R}(I,J)$. Therefore, we may assume that $\operatorname{cx}_{R}(k)$ is finite. By Proposition 2.7.(2), this implies that $R$ is a complete intersection. Hence, Propositions 2.9.(1) and 2.7.(1) yield $\operatorname{cx}_{R}(k)\leqslant\operatorname{cx}_{R}(I,J)\leqslant\min\{\operatorname{cx}_{R}(I),\operatorname{cx}_{R}(J)\}\leqslant\operatorname{cx}_{R}(k)$. Consequently, $\operatorname{cx}_{R}(I,J)=\operatorname{cx}_{R}(k)$. Thus, the equivalence of (1) and (2) follows from Proposition 2.7.(2). Since $\operatorname{curv}_{R}(k)\leqslant\operatorname{curv}_{R}(I,J)$, the equivalence of (1) and (3) can be observed from Propositions 2.7.(2) and 2.9.(1). ∎
4. Some examples
Here we provide two examples computing complexity and curvature for pairs of Burch ideals to complement our results.
Example 4.1.
Let $(R,\mathfrak{m},k)$ be a Noetherian local ring such that $\mathfrak{m}\neq 0$ and $\mathfrak{m}^{2}=0$. Set $b:=\mu(\mathfrak{m})$. For example, $R$ can be $k[x_{1},\ldots,x_{b}]/(x_{i}x_{j}:1\leqslant i\leqslant j\leqslant b)$, where $x_{1},\ldots,x_{b}$ are indeterminates over a field $k$. Note that $\mathfrak{m}$ is a Burch ideal of $R$. Moreover,
$\operatorname{tcx}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{cx}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{cx}_{R}(k)=\begin{cases}1&\mbox{if $b=1$,}\\ \infty&\mbox{if $b\geqslant 2$,}\end{cases}$
and $\operatorname{tcurv}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{curv}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{curv}_{R}(k)=b$. Clearly, $R$ is complete intersection if and only if $b=1$ if and only if $R$ is a hypersurface.
Proof.
Note that $\mathfrak{m}\cong k^{\oplus b}$. Then $\beta_{n}^{R}(k)=\beta_{n-1}^{R}(\mathfrak{m})=b\beta_{n-1}^{R}(k)$ for all $n\geqslant 1$. Thus $\beta_{n}^{R}(k)=b\beta_{n-1}^{R}(k)=b^{2}\beta_{n-2}^{R}(k)=\cdots=b^{n}\beta_{0}^{R}(k)=b^{n}$ for all $n$. Now $\operatorname{Tor}_{n}^{R}(\mathfrak{m},\mathfrak{m})\cong\operatorname{Tor}_{n}^{R}(k^{\oplus b},k^{\oplus b})\cong\bigoplus_{b}\bigoplus_{b}\operatorname{Tor}_{n}^{R}(k,k)$. Similarly, $\operatorname{Ext}^{n}_{R}(\mathfrak{m},\mathfrak{m})\cong\bigoplus_{b}\bigoplus_{b}\operatorname{Ext}^{n}_{R}(k,k)$. So $\mu_{R}(\operatorname{Tor}_{n}^{R}(\mathfrak{m},\mathfrak{m}))=\mu_{R}(\operatorname{Ext}^{n}_{R}(\mathfrak{m},\mathfrak{m}))=b^{2}\beta_{n}^{R}(k)=b^{n+2}$ for all $n$. Hence, computing complexity and curvature, the desired equalities follow. ∎
Our next example shows that the $\mathfrak{m}$-primary condition on $I$ is not a necessary condition in Corollary 1.5. This leads to 4.3.
Example 4.2.
Fix an integer $b\geqslant 1$. Let $R:=k[x_{1},\ldots,x_{b}]/(x_{i}x_{j}:1\leqslant i<j\leqslant b)$, where $x_{1},\ldots,x_{b}$ are indeterminates over a field $k$. Set $\mathfrak{m}:=(x_{1},\ldots,x_{b})$. Let $\ell,m\geqslant 1$ be two integers. Then $\mathfrak{m}^{\ell}$ and $(x_{i}^{\ell+1})=x_{i}^{\ell}\mathfrak{m}$ (for $1\leqslant i\leqslant b$) are Burch ideals, however $(x_{i}^{\ell})$ are not $\mathfrak{m}$-primary. For $1\leqslant i,j\leqslant b$, we have
$\operatorname{tcx}_{R}((x_{i}^{\ell}),(x_{j}^{m}))=\operatorname{tcx}_{R}((x_{i}^{\ell}),\mathfrak{m}^{m})=\operatorname{tcx}_{R}(\mathfrak{m}^{\ell},\mathfrak{m}^{m})=\operatorname{cx}_{R}((x_{i}^{\ell}),(x_{j}^{m}))$
$=\operatorname{cx}_{R}((x_{i}^{\ell}),\mathfrak{m}^{m})=\operatorname{cx}_{R}(\mathfrak{m}^{\ell},(x_{j}^{m}))=\operatorname{cx}_{R}(\mathfrak{m}^{\ell},\mathfrak{m}^{m})=\operatorname{cx}_{R}(k)=\begin{cases}b-1&\mbox{if $1\leqslant b\leqslant 2$,}\\ \infty&\mbox{if $b\geqslant 3$,}\end{cases}$
and
$\operatorname{tcurv}_{R}((x_{i}^{\ell}),(x_{j}^{m}))=\operatorname{tcurv}_{R}((x_{i}^{\ell}),\mathfrak{m}^{m})=\operatorname{tcurv}_{R}(\mathfrak{m}^{\ell},\mathfrak{m}^{m})=\operatorname{curv}_{R}((x_{i}^{\ell}),(x_{j}^{m}))$
$=\operatorname{curv}_{R}((x_{i}^{\ell}),\mathfrak{m}^{m})=\operatorname{curv}_{R}(\mathfrak{m}^{\ell},(x_{j}^{m}))=\operatorname{curv}_{R}(\mathfrak{m}^{\ell},\mathfrak{m}^{m})=\operatorname{curv}_{R}(k)=b-1.$
Proof.
Fix $1\leqslant i,j\leqslant b$. Observe that $\operatorname{ann}_{R}(x_{i}^{\ell})=\operatorname{ann}_{R}(x_{i})$. Therefore, $(x_{i}^{\ell})\cong R/\operatorname{ann}_{R}(x_{i}^{\ell})\cong R/\operatorname{ann}_{R}(x_{i})\cong(x_{i})$. Consequently, $\mathfrak{m}^{\ell}\cong(x_{1}^{\ell})\oplus\dots\oplus(x_{b}^{\ell})\cong(x_{1})\oplus\dots\oplus(x_{b})\cong\mathfrak{m}$. So we need to compute complexity and curvature involving $(x_{i})$, $(x_{j})$ and $\mathfrak{m}$.
Since $\mathfrak{m}\cong(x_{1})\oplus\dots\oplus(x_{b})$, it follows that
(4.1) $\bigoplus_{r=1}^{b}\operatorname{Ext}^{n}_{R}((x_{r}),\mathfrak{m})\cong\operatorname{Ext}^{n}_{R}(\mathfrak{m},\mathfrak{m})\cong\bigoplus_{s=1}^{b}\operatorname{Ext}^{n}_{R}(\mathfrak{m},(x_{s}))\cong\bigoplus_{r,s=1}^{b}\operatorname{Ext}^{n}_{R}((x_{r}),(x_{s})).$
Hence, by Lemma 2.2, $\operatorname{cx}_{R}(\mathfrak{m},\mathfrak{m})=\sup\{\operatorname{cx}_{R}((x_{r}),\mathfrak{m}):1\leqslant r\leqslant b\}=\operatorname{cx}_{R}((x_{i}),\mathfrak{m})$ by symmetry of the variables. Similarly, $\operatorname{cx}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{cx}_{R}(\mathfrak{m},(x_{j}))$. Moreover, the isomorphisms (4.1) yield that $\operatorname{cx}_{R}(\mathfrak{m},\mathfrak{m})=\sup\{\operatorname{cx}_{R}((x_{r}),(x_{s})):1\leqslant r,s\leqslant b\}$. By symmetry of the variables, note that $\operatorname{cx}_{R}((x_{r}),(x_{s}))$, for $1\leqslant r,s\leqslant b$ with $r\neq s$, are equal. Also $\operatorname{cx}_{R}((x_{r}),(x_{r}))$, for $1\leqslant r\leqslant b$, are equal. Since $\Omega_{R}^{1}(x_{r})=\bigoplus_{1\leqslant s\leqslant b,s\neq r}(x_{s})$, in view of Lemma 2.2, $\operatorname{cx}_{R}((x_{r}),(x_{r}))=\operatorname{cx}_{R}(\Omega_{R}^{1}(x_{r}),(x_{r}))=\sup\{\operatorname{cx}_{R}((x_{s}),(x_{r})):1\leqslant s\leqslant b,s\neq r\}=\operatorname{cx}_{R}((x_{s}),(x_{r}))$ for every $s\neq r$. Thus, $\operatorname{cx}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{cx}_{R}((x_{i}),(x_{j}))$. Combining all these equalities, $\operatorname{cx}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{cx}_{R}((x_{i}),\mathfrak{m})=\operatorname{cx}_{R}(\mathfrak{m},(x_{j}))=\operatorname{cx}_{R}((x_{i}),(x_{j}))$. Similarly, one obtains the corresponding equalities for curvature, Tor-complexity and Tor-curvature.
Since $\mathfrak{m}=\Omega_{R}^{1}(k)$ and $\mathfrak{m}$ is a Burch ideal, in view of Proposition 2.20, $\operatorname{cx}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{cx}_{R}(k,\mathfrak{m})=\operatorname{inj\,cx}_{R}(\mathfrak{m})=\operatorname{cx}_{R}(k)$ and $\operatorname{tcx}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{tcx}_{R}(k,\mathfrak{m})=\operatorname{cx}_{R}(\mathfrak{m})=\operatorname{cx}_{R}(k)$. Similarly, $\operatorname{curv}_{R}(\mathfrak{m},\mathfrak{m})=\operatorname{curv}_{R}(k)=\operatorname{tcurv}_{R}(\mathfrak{m},\mathfrak{m})$, which is same as $\operatorname{curv}_{R}(\mathfrak{m})$. So, in order to prove the desired equalities, it is enough to compute $\operatorname{cx}_{R}(\mathfrak{m})$ and $\operatorname{curv}_{R}(\mathfrak{m})$.
Since $\Omega_{R}^{1}(x_{r})=\bigoplus_{1\leqslant s\leqslant b,s\neq r}(x_{s})$, it follows that $\beta_{1}^{R}((x_{r}))=\mu(\Omega_{R}^{1}(x_{r}))=b-1$ for $1\leqslant r\leqslant b$. Hence
$\beta_{2}^{R}(x_{r})=\beta_{1}^{R}(\Omega_{R}^{1}(x_{r}))=\beta_{1}^{R}(\bigoplus_{1\leqslant s\leqslant b,s\neq r}(x_{s}))=\sum_{1\leqslant s\leqslant b,s\neq r}\beta_{1}^{R}((x_{s}))=(b-1)^{2}$.
Thus, using induction, for each $n\geqslant 1$, we obtain that $\beta_{n}^{R}(x_{r})=\beta_{n-1}^{R}(\Omega^{1}_{R}(x_{r}))=(b-1)^{n}$ for $1\leqslant r\leqslant b$. Therefore $\beta_{n}^{R}(\mathfrak{m})=\beta_{n}^{R}\big((x_{1})\oplus\dots\oplus(x_{b})\big)=\sum_{r=1}^{b}\beta_{n}^{R}(x_{r})=b(b-1)^{n}$ for all $n\geqslant 1$. So
$\operatorname{cx}_{R}(\mathfrak{m})=\begin{cases}0&\mbox{if $b=1$,}\\ 1&\mbox{if $b=2$,}\\ \infty&\mbox{if $b\geqslant 3$,}\end{cases}$
and $\operatorname{curv}_{R}(\mathfrak{m})=b-1$. Thus the required equalities follow. ∎
We close this article by presenting the following natural questions.
Question 4.3.
Does Corollary 1.5 hold even if $I$ is not $\mathfrak{m}$-primary?
Question 4.4.
Does Corollary 1.6 hold even if $\widehat{R}$ is not locally Gorenstein on punctured spectrum?